Integrals Giving Inverse Trigonometric Functions


∫(e^x dx)/√(1-e^(2x))


1
Expert's answer
2021-04-15T07:40:03-0400

Put "e^x=u", then "x=ln u, dx = \frac{du}{u}". Also, "e^{2x}=(e^x)^2=u^2". Hence,


"\intop \frac{e^x dx}{\sqrt{1-e^{2x}}}= \intop \frac{u\cdot \frac{du}{u}}{\sqrt{1-u^2}}=\intop \frac{du}{\sqrt{1-u^2}}="

"=sin^{-1}u +C=sin^{-1}e^x +C"

Answer: "sin^{-1}e^x +C"